OpenStudy (anonymous):

Profile as a power of 3 a) 1/81 b) 9/27

OpenStudy (anonymous):

a) 3^-4 b) 3.3^-3

OpenStudy (anonymous):

how did u do that?

OpenStudy (shane_b):

a) \[\frac{1}{81}=\frac{1}{3^4}=3^{-4}\]b) \[\frac{9}{27}=\frac{3^2}{3^3}=3^{-1}\]

OpenStudy (shane_b):

Does that help?

OpenStudy (anonymous):

ye thnx